OpenStudy (anonymous):

When solving an exponential equation, why is it necessary to use logarithms?

OpenStudy (dumbcow):

because logs are defined as the inverse function to the exponential function: e^x = 5 --> x = ln(5) e^(ln5) = 5

OpenStudy (dumbcow):

same reason we use square roots to solve quadratic equations

OpenStudy (anonymous):

Thank you